OpenStudy (anonymous):

Translate to an equation, then solve. Thirteen less than six times a number is equal to 11 more than four times the number . What is the number ?

OpenStudy (anonymous):

say that no is "x" then 13-6x=11+4x solving for x 4x+6x=13-11 =>10x=2 =>x=1/5

OpenStudy (anonymous):

that cannot be a number rite ? only fraction ?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yeeeppp..have you choices..??